Problem 10.107 Constants 1 Periodic Table Assume that an exhaled breath of air consists of 747%N2 15 4 %02.36%CO2 and 63%water vapor
media%2F673%2F673675bb-3af1-431a-8318-2e
media%2F3dd%2F3dd1d0e4-0d31-4078-a6b1-8c
media%2F39f%2F39f6cca9-e0f2-4c0b-949a-da
Answer #1

Part-A
Total pressure = P = 0.990 atm
molefraction of N2= 74.7/100 =0.747
Partial pressure of N2= Total pressure x molefraction of N2
Partial pressure = 0.990x0747 =0.73953
partail pressure of N2 =0.739 atm

Part-B
Total pressure = 0.990 atm
Molefraction of O2= 15.4/100 = 0.154 atm
partial pressure of O2= 0.990x0.154 =0.15246
partial pressure of O2 = 0.153 atm

Part-C
Total pressure = 0.990 atm
molefraction of CO2 = 3.6/100 = 0.036atm
Partail pressure of CO2 = 0.990x0.036 =0.03564
Partial pressure of CO2 = 0.0356 atm

Part-D
Total pressure = 0.990 atm
molefraction of H2O = 6.3/100 = 0.063
partial presssure of H2O = 0.990x0.063 =0.06237
partial pressure of H2O = 0.0624 atm.

Part-E
Partial pressure of CO2 = 0.0356 atm
volume= 465ml=0.465L
T= 37C=37+273= 310K
R=0.0821L-arm/mole-K
PV=nRT
n=PV/RT = 0.0356x0.465/0.0821x310 =0.00065 moles
numberr of moles of CO2= 0.00065 moles